OpenStudy (anonymous):

CAN SOMEONE PLEASE PLEASE HELP ME!!!! I just took my semester exam (which I cannot retake) and I got this question wrong (I was so confused) will someone for the love of god please help me understand this question Simplify (4sqrt6)/(5sqrt6) A. 6^1/5 B. 6^9/20 C 6^1/20 D 6^5/4 ITS NOT D

OpenStudy (anonymous):

@Tom_Boy_Rebel @Vocaloid @MissNoName

OpenStudy (anonymous):

I WILL MEDAL AND FAN!

RhondaSommer (rhondasommer):

give me just a minute and i can help you

OpenStudy (anonymous):

OH THANK THE LORD

OpenStudy (anonymous):

6^1/5 or 6^(1/5)

OpenStudy (anonymous):

Is this the expression? \[\frac{ 4\sqrt{6} }{ 5\sqrt{6} }\]

RhondaSommer (rhondasommer):

okay so each the denominator and the numerator has √6 corrrect?

OpenStudy (anonymous):

Yes

RhondaSommer (rhondasommer):

then they cancel out :)

RhondaSommer (rhondasommer):

and you are left with what/what

OpenStudy (anonymous):

Yea so then u would be left with 4/5 I already know that I even did it with an algebra calculator but my brain doesn't want to comprehend which answer it is I apologize for my stupidity

RhondaSommer (rhondasommer):

? you arent stupid. did you answer with 4/5 or 0.8 and get it wrong?

OpenStudy (anonymous):

It's not an answer choice look up I putt he answer choices

OpenStudy (anonymous):

Your answers were (just to check): A. \[6^{\frac{ 1 }{ 5 }}\] B. \[6^{\frac{ 9 }{ 20 }}\] C. \[6^{\frac{ 1 }{ 20 }}\] D. \[6^{\frac{ 5 }{ 4 }}\]

OpenStudy (anonymous):

Yes correct I know it's not D

RhondaSommer (rhondasommer):

thats interesting bc none of those answers make sense...

OpenStudy (anonymous):

I know I literally started crying

OpenStudy (anonymous):

I put all of those into a calculator and none are equal to 4/5. Did you ask your teacher?

RhondaSommer (rhondasommer):

your teacher is incorrect in this situation. sorry bruh :/ you got it right tho!

OpenStudy (anonymous):

Will I cant bc its winter break and plus it was a semester exam which means I'm done with the semester I can go back

OpenStudy (anonymous):

Can't *

RhondaSommer (rhondasommer):

sorry bruh :/ maybe you can request a retake bc it was a teacher typo not your misunderstanding.

OpenStudy (anonymous):

Your teacher should be able to change your overall grade even after the semester has ended. I wouldn't be too worried about it if he/she made the mistake. After break just explain what happened and I'm sure it will work out.

OpenStudy (anonymous):

Will I don't believe he can really reset it and I happy with my grade (93) but I just wanted to understand it for the eoc

OpenStudy (anonymous):

Oh well thank you for trying

OpenStudy (anonymous):

Can you post a pic of the question?

OpenStudy (anonymous):

Um I'm on my phone and I don't exactly know how

OpenStudy (anonymous):

Thank you all for attempting to help:)

OpenStudy (anonymous):

Should I close the question now?